Lisa consumes only pizzas and burritos. In equilibrium, her marginal utility of pizza is 32 and her marginal utility of a burrito is 24. The price of a pizza is $4.

*Round your answer to 2 decimal places.

The price of a burrito is $____.
Answer #1

ANSWER:

In equilibrium Marginal utility of pizza divided by price of pizza will be equal to marginal utility of burritos divided by price of burrito.

let the price of burrito be x.

therefore, 32 / 4 = 24 / x

8 = 24 / x

x = 24 / 8 = 3

therefore the price of burrito is $3.
Answer #2

 In equilibrium, the ratio of the marginal utility of a pizza divided by the price of a pizza must equal the marginal utility of a burrito divided by the price of a burrito. Thus, the price of a burrito must be $2.